OpenStudy (anonymous):

Please help :) and explaining it would be awesome! Find the indicated value of the function. f(x)=sqrt (x+5) -x-3 f ( sqrt (2)+3 ) = ?

OpenStudy (anonymous):

\[f(\sqrt{2}+3)=\sqrt{(\sqrt{2}+3)+5}-(\sqrt{2}+3)-3\]

OpenStudy (precal):

\[\sqrt{\sqrt{2}+5}-\sqrt{2}-3-3\] \[\sqrt{\sqrt{2}+5}-\sqrt{2}-6\]

OpenStudy (anonymous):

thanks! yet, these are not coming up as right answers....is there anyway they can be simplified more maybe?

OpenStudy (precal):

as decimals

OpenStudy (anonymous):

it wouldn't be that...

OpenStudy (precal):

do you have choices

OpenStudy (anonymous):

no i dont :(

OpenStudy (precal):

read the directions

OpenStudy (precal):

did you type it in with the square roots? sometimes it is your syntax (use parenthesis)

OpenStudy (anonymous):

ha thats all they gave me, its okay thanks though!

OpenStudy (precal):

ok sorry that is the limitations of technology